# BWTP model for hadrons $p_{\rm T}$-spectra and elliptic flow $v_{2}$ in AA collisions at the LHC

### v1

@@ Abstract
A generalization of the phenomenological BWTP model, recently proposed by the author for the hadrons transverse momentum ($p_{\textrm{T}}$) spectra measured at the LHC, is developed to also describe the hadrons $p_{\textrm{T}}$-differential elliptic flow coefficients $v_{2}$ in identical nuclei (AA) collisions of different centralities. The model describes well the available data on $p_{\textrm{T}}$-spectra and $v_{2}$ for any $p_{\textrm{T}}$ of various particles, from pions to charmonia, in Pb–Pb at $\sqrt{s_{\mathrm{NN}}}=$ 2.76 and 5.02 TeV, and in Xe–Xe at $\sqrt{s_{\mathrm{NN}}}=$ 5.44 TeV. Also, predictions for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$ 5.36 TeV are given. While the model is mainly targeting the LHC energies, it also works at much lower RHIC energies.
@@ I Introduction
Recently, the BWTP model (BWTPM) was proposed, which describes quite accurately the midrapidity $p_{\textrm{T}}$-spectra of various hadrons in the whole $p_{\textrm{T}}$ range measured in pp and Pb–Pb collisions at the LHC [1]. These $p_{\textrm{T}}$-spectra are integrated over the hadron azimuthal angle $\varphi$. They depend strongly on the collision centrality and centre-of-mass energy per nucleon-nucleon pair $\sqrt{s_{\mathrm{NN}}}$. BWTPM has three terms including a standard blast-wave model (BWM) term [2] to account for the hydrodynamical radial flow and kinetic freeze-out effects, a term in the Tsallis (Ts) distribution form [3, 4], responsible for the contribution from the resonance decays, and a power-law (Po) term describing the QCD hard processes. As shown in Ref. [1], the model parameters depend generally on the collision system (pp or Pb–Pb), but some of them are independent of the $\sqrt{s_{\mathrm{NN}}}$ or the particle type or the collision centrality.
In the present paper, a generalization of the BWTPM is developed to describe, in addition to the $p_{\textrm{T}}$-spectra, also the midrapidity elliptic flow coefficients $v_{2}$ in identical nuclei (AA) collisions. The $v_{2}$ characterizes the azimuthal asymmetry of hadron momentum distribution and depends on hadron type, $p_{\textrm{T}}$, collision system, and centrality. The asymmetry is driven mainly by an almond shape (in the plane transverse to the collision axis) of the hadronic medium (quark-gluon plasma) created in non-central collisions. Due to this shape, the in-medium pressure gradient and radial flow velocity become anisotropical resulting in the anisotropy of the hadron momentum spectra [5]. Following to Ref. [6], we use the generalized BWM with transversally anisotropic freeze-out hyper-surface of elliptical shape, instead of the isotropic circular one considered in Refs. [1, 2]. For the hadronic medium radial flow velocity we assume similar azimuthal dependence as in Refs. [6, 7]. Also, the other two therms of BWTPM are modified by including a simple dependence on $\varphi$. Having redefined so the model, the hadron $v_{2}$ can be determined as the second coefficient in the Fourier series of its invariant yield $\textrm{d}^{3}N/(p_{\textrm{T}}\textrm{d}p_{\textrm{T}}\textrm{d}y\textrm{d}\varphi)$ [6, 7, 8, 9, 10]
@@
Next objective of the present paper is to reduce strongly the number of the BWTPM parameters by parameterizing their dependence on the collision system, centrality, energy, and on the particle type and $p_{\textrm{T}}$ using simple formulae. As a measure of collision centrality, the corresponding charged-particle multiplicity density at midrapidity ($\eta$ is pseudo-rapidity)
@@
is used which serves as a good scaling variable for the comparison of hadron production mechanisms in different collision systems (see, e.g., the ALICE review paper [11]). These parameterizations increase significantly the predictive power of the model. In addition, we introduce another simplification of the BWTPM. While in Ref. [1] it has two different functional forms for $p_{\textrm{T}}\leq 40$ GeV and $p_{\textrm{T}}>40$ GeV, here it has the same form for any $p_{\textrm{T}}$.
The values of the BWTPM parameters are obtained then from a simultaneous fit of the combined data on different hadron $p_{\textrm{T}}$-spectra and $v_{2}$ measured mostly by the ALICE experiment in Pb–Pb and Xe–Xe collisions at the LHC. In addition to the hadrons considered in Ref. [1] ($\pi^{\pm}$, $\pi^{0}$, $\mathrm{K}^{\pm}$, ${\mathrm{K}}^{0}_{\mathrm{S}}$, $\mathrm{p}$, $\mathrm{\overline{p}}$, ${\mathrm{K}}^{*0}$, $\overline{\mathrm{K}^{*0}\mkern-15.0mu}$ , $\phi$, $\Lambda$, $\overline{\Lambda}$, $\Xi^{-}$, $\overline{\Xi^{+}\mkern-12.0mu}$ , $\Omega^{-}$, $\overline{\Omega^{+}\mkern-12.0mu}$ , ${\mathrm{D}}^{0}$, $\mathrm{D}^{\pm}$, $\mathrm{J}/\psi$, and unidentified charged-particles), the $\eta$, $\rho^{0}$ mesons and $\Lambda_{c}^{+}$$\overline{\Lambda^{-}\mkern-12.0mu}_{c}$ baryons are included too. We assume that a particle and its antiparticle have the same $p_{\textrm{T}}$-spectrum and $v_{2}$. While in the LHC energy domain, which is the main target of our study, this assumption is perfectly valid, it could be violated at much lower energies. To test our model at such energies, we included in the simultaneous fit also some limited data on $p_{\textrm{T}}$-spectra and $v_{2}$ measured at the RHIC.
Note that the previous similar simultaneous fits of the LHC data on hadrons $p_{\textrm{T}}$-spectra and $v_{2}$ were done mainly at low $p_{\textrm{T}}$ ($p_{\textrm{T}}\lesssim 3$ GeV/c[7, 11, 12] using various generalizations of the BWM. Contrary, our model, due to its additional power-law terms, is able to describe much larger data set in the whole available $p_{\textrm{T}}$-region.
The paper is organized as follows: Sec. II gives details of the generalized BWTPM. Sec. III presents a parameterization for the $M_{\rm ch}$ data measured in various AA collisions. Sec. IV describes parameterizations for the model parameters dependence on the AA collision characteristics and the particle type and $p_{\textrm{T}}$. In Sec. V the model free parameters are determined from the simultaneous fit of all the used data. In Sec. VI the results and predictions are discussed. Conclusions are given in the last section.
@@ II Description of BWTPM
Present model consists of three terms which generalize the similar terms of Ref. [1]. The first (BW) term corresponds to the generalized BWM [6] and two others (named Ts and Po as in Ref. [1]) include additionally the second Fourier coefficients $v_{2t}(p_{\textrm{T}})$ and $v_{2p}(p_{\textrm{T}})$, respectively, to account for the elliptic flow. Thus, for a particle $\varphi$-dependent $p_{\textrm{T}}$-spectrum in AA collisions at midrapidity, where it can be considered as independent of the rapidity $y$, we propose the following model
@@
$\rho$ is the flow rapidity with the second order anisotropy $\rho_{2}$ and profile exponent $k$, $V=\pi R_{x}R_{y}\tau_{f}$ is the volume of this medium with transverse sizes $R_{x}$ and $R_{y}$, $\tau_{f}$ and $T$ are its proper life-time and temperature, $K_{1}$ is a modified Bessel function of the second kind. Note that $\tanh{\rho_{0}}$ is equal to $\beta_{\textrm{s}}$ used in Ref. [1]. The global normalization parameter $f_{N}$ is close to unity at LHC energies but could depend on the normalization conditions of the used data obtained in different experiments. Its energy dependence is significant at lower RHIC energies and will be specified in Sec. IV. Parameter $f_{S}$ is equal unity (as in Ref [1]) for ordinary hadrons consisting of up and down quarks only. But for hadrons containing strange or heavier quarks it is greater unity and quantifies the suppression of their yields depending on the hadron type, $p_{\textrm{T}}$, and collision centrality. It describes the observed enhancement of the relative yields of such hadrons with respect to ordinary ones when going from peripheral to central collisions (i.e., from small to large $M_{\rm ch}$[11]. The explicit form of $f_{S}$ is specified in Sec. IV.
@@
are explained in Refs. [6, 7]. In our model, three parameters $R_{x},R_{y}$, and $\tau_{f}$ enter into two independent combinations via parameters $V$ and $r_{xy}\leq 1$.
The important quantities for the elliptic flow $v_{2}$ are $\rho_{2},r_{xy}$ [6, 7], $k$, and functions $v_{2t}(p_{\textrm{T}})$ and $v_{2p}(p_{\textrm{T}})$ in Eq. (II), though the $T$ and $\rho_{0}$ also influence the $p_{\textrm{T}}$-dependence of $v_{2}$ [6]. From hydrodynamic models and experimental data it follows an approximate linear relation between $v_{2}$ and the eccentricity $\varepsilon_{2}$ of the collision geometry [13, 14, 15, 16], which can be defined in the present model as
@@
Moreover, in the generalized BWM [6] with $\rho_{2}=0$ one gets $v_{2}=\varepsilon_{2}/2$. We have checked by fitting the $v_{2}$ data that parameter $\rho_{2}$ as well as the functions $v_{2t}(p_{\textrm{T}})$ and $v_{2p}(p_{\textrm{T}})$ have similar to $\varepsilon_{2}$ centrality dependence. For the model simplicity, we assume that they are proportional to $\varepsilon_{2}$. Their explicit forms will be specified in Sec. IV. All other parameters in Eq. (II) have the same meaning as in Ref. [1]. Note that parameters $f_{BW},f_{1},f_{2},c_{1}$, and $c_{2}$ are the same for any AA collision system and depend only on the particle type.
@@
Performing similar integration in Eq. (1) one finds (see also Ref. [7])
@@
The $I_{0}$ and $I_{2}$ in Eqs. (II) and (II) are modified Bessel functions of the first kind. So, we found Eqs. (8), (II), and (II) for hadrons $p_{\textrm{T}}$-spectrum and elliptic flow, respectively. They can be used to describe the corresponding data for different identified particles. There are also similar data sets for the unidentified charged particles which are a mixture of the $\pi^{\pm},\mathrm{K}^{\pm},\mathrm{p},\mathrm{\overline{p}}$, and charged hyperons. These data have usually much larger $p_{\textrm{T}}$ reach than the ones for the identified particles. Charged-particle $p_{\textrm{T}}$-spectra can be defined, assuming same spectrum for a particle and its anti-particle, by the equations [1]
@@
Here, the functions $F_{\varphi,i}$ and $F_{i}$ are defined by Eqs. (II) and (II), respectively. The factors 2 account for the positive and negative particles and factors $p_{\textrm{T}}/m_{\textrm{T}}$ account for the change from the rapidity to pseudo-rapidity at midrapidity ($d\eta/dy\approx p_{\textrm{T}}/m_{\textrm{T}}$). The last terms in Eqs. (11) and (12) describe approximately the small contribution of hyperons via the proton contribution, scaled by parameter $f_{hyp}$. Using Eqs. (11) and (1) and performing integration by $\varphi$ one can find easily the analogue of Eq. (II) for the unidentified charged particles. Note that the Eqs. (8), (II) and (12) for the $p_{\textrm{T}}$-spectra are similar to the corresponding ones in Ref. [1] if $f_{N}=f_{S}=1$ and $\rho_{2}=0$.
@@
We have checked that the presented generalized BWTPM is able to fit the LHC data on midrapidity hadron $p_{\textrm{T}}$-spectra and $v_{2}$ in Pb–Pb and Xe–Xe collisions with a good accuracy, similar to one achieved in Ref. [1]. However, as in Ref. [1], the number of the model free parameters is too big, which reduces the model’s predictive power. It is mostly due to the following ten centrality-dependent parameters $V,V_{1},V_{2},T,k,e_{2},n,f_{S},r_{xy},$ and $\rho_{0}$ which generally can depend also on the $p_{\textrm{T}}$, collision system and $\sqrt{s_{\mathrm{NN}}}$ ($f_{S}$ depends on particle species too). Typical number of the ALICE centrality classes for Pb–Pb data at $\sqrt{s_{\mathrm{NN}}}=$ 2.76 TeV [17] and 5.02 TeV [18] as well as for Xe–Xe data at $\sqrt{s_{\mathrm{NN}}}=$ 5.44 TeV  [19] is about 10 in each case. So, to describe these combined data, one needs to fit $10(10+10+10)=300$ independent values for the mentioned ten parameters. Taking into account also the RHIC data included in the present study, this number of independent values becomes more than 600. To reduce the number of free parameters strongly, we will parameterize these ten (and other) parameters by using the charged-particle multiplicities $M_{\rm ch}$ which depend on the collision system, centrality and $\sqrt{s_{\mathrm{NN}}}$, and are well known experimentally for various AA collisions  [17, 18, 19, 20, 21, 22]. The key assumption of the present paper is that all the centrality dependence in BWTPM can be expressed via the $M_{\rm ch}$ only. It is motivated by the experimental observation that any hadron species has similar $p_{\textrm{T}}$-spectra in different AA collisions with the same $M_{\rm ch}$ and $\sqrt{s_{\mathrm{NN}}}$ (see an example of such a similarity for ${\mathrm{K}}^{*0}$ meson $p_{\textrm{T}}$-spectra in Ref. [23]). To have a more general and predictive approach, we will use not the experimental values of $M_{\rm ch}$ (which can be missing for some centralities), but their accurate parameterization, valid for any AA collision centrality and energy. This parameterization is presented in the next section.
III parameterization of $M_{\rm ch}$
@@
The mean value of charged-particle multiplicity density at midrapidity $M_{\rm ch}$ (defined in Eq. (2)) is one of the main characteristics of nucleus-nucleus collisions. Its dependence on the collision system, centrality and $\sqrt{s_{\mathrm{NN}}}$ was measured for different AA collisions at the LHC [17, 18, 19, 20] and RHIC [21, 22] energies. Centrality is commonly expressed in percentiles or fractions $x_{c}$ of the total nucleus-nucleus cross section [11] (centrality in % is equal to $100x_{c}$). Here we propose a simple parameterization for the $M_{\rm ch}$ data [17, 18, 19, 20, 21, 22] versus atomic mass number A, collision centrality fraction $x_{c}$ and energy $\sqrt{s_{\mathrm{NN}}}$. It relates $M_{\rm ch}$ with the similar quantity $M^{\rm pp}_{\rm ch}\equiv\langle\textrm{d}N^{\rm pp}_{\rm ch}/\textrm{d}\eta\rangle_{\eta=0}$ of the inelastic pp collisions at energy $\sqrt{s}=\sqrt{s_{\mathrm{NN}}}$. For $M^{\rm pp}_{\rm ch}$ we use the data measured at the LHC [24, 25, 26] as well as at lower energies [21, 27] and parameterize its energy dependence. The $M_{\rm ch}$ and $M^{\rm pp}_{\rm ch}$ data were obtained by averaging the respective multiplicity distribution on pseudo-rapidity in the range $|\eta|<0.5$ [17, 18, 19, 20, 22, 24, 25, 26] or $|\eta|<1$ [21, 27]. Within the data uncertainties both ranges give similar results [26]. Parameters of the parameterization are obtained via a simultaneous fit of the $M_{\rm ch}$ and $M^{\rm pp}_{\rm ch}$ data using the combined fit technique of the ROOT framework [28], which gives $\chi^{2}/NDF=54.1/155$. The fit results are shown in Figures 1 and 2 and the parameterization has the form
@@
where $\alpha=\log(\frac{17.4}{\sqrt{s_{\mathrm{NN}}}}),\beta=(\frac{0.0027}{\sqrt{s_{\mathrm{NN}}}})^{0.054},\delta=\textrm{A}^{0.063},\gamma=1+1.22x_{c}^{0.47},y=\textrm{A}(1-x_{c})$, and $e_{0}=0.005$ TeV (everywhere in the present paper, if it is not specified otherwise, $\sqrt{s_{\mathrm{NN}}}$ is defined in units of TeV). The bottom panel of Fig. 1 demonstrates that Eq. (13) describes most of the data-points with better than 10% accuracy. However, it underestimates the ALICE recent data for most peripheral Pb–Pb collisions at 5.36 TeV [20]. This could be due to the new centrality calibration [20], since all other ALICE data are described well. One can expect that Eq. (13) would give accurate values of $M_{\rm ch}$ also for new AA collisions. Result for oxygen–oxygen (OO) collisions, carried out at the LHC in July 2025, is also shown in Fig. 1. Note that for $\textrm{A}=1$ or for most peripheral collisions, when $x_{c}\to 1$, the right-hand side of Eq. (13) equals unity. For most central collisions, when $x_{c}\to 0$, we obtain a very simple formula for the maximal values of $M_{\rm ch}$ versus A and $\sqrt{s_{\mathrm{NN}}}$
$\frac{M_{\rm ch}^{max}}{M^{\rm pp}_{\rm ch}}=1+(\textrm{A}-1)^{1+0.00013\textrm{A}}(\frac{\sqrt{s_{\mathrm{NN}}}}{0.0027})^{0.1134}.$ (15)
For example, $M_{\rm ch}^{max}=$ 2404 (2457) for Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}=$5.02 (5.36) TeV.

### v3

# Blast-wave-Tsallis-power model for $p_{\rm T}$-spectra and elliptic flow $v_{2}$ of hadrons in collisions of identical nuclei at energies available at the Large Hadron Collider

@@ Abstract
A generalization of the phenomenological blast-wave-Tsallis-power model, recently proposed by the author for the hadrons transverse momentum ($p_{\textrm{T}}$) spectra measured at the LHC, is developed to also describe the hadrons $p_{\textrm{T}}$-differential elliptic flow coefficients $v_{2}$ in identical nuclei collisions of different centralities. The model describes well the available data on $p_{\textrm{T}}$-spectra and $v_{2}$ for any $p_{\textrm{T}}$ of various particles, from pions to charmonia, in Pb–Pb at $\sqrt{s_{\mathrm{NN}}}=$ 2.76 and 5.02 TeV, and in Xe–Xe at $\sqrt{s_{\mathrm{NN}}}=$ 5.44 TeV. Also, predictions for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$ 5.36 TeV are given. While the model is mainly targeting the LHC energies, it also works at much lower RHIC energies.
@@ I Introduction
Recently, the phenomenological blast-wave-Tsallis-power model (BWTPM) was proposed, which describes quite accurately the midrapidity $p_{\textrm{T}}$-spectra of various hadrons in the whole $p_{\textrm{T}}$ range measured in pp and Pb–Pb collisions at the LHC [1]. These $p_{\textrm{T}}$-spectra are integrated over the hadron azimuthal angle $\varphi$. They depend strongly on the collision centrality and center-of-mass energy per nucleon-nucleon pair $\sqrt{s_{\mathrm{NN}}}$. BWTPM has three terms including a standard blast-wave model (BWM) term [2] to account for the hydrodynamic radial flow and kinetic freeze-out effects, a term in the Tsallis (Ts) distribution form [3, 4], responsible for the contribution from the resonance decays, and a power-law (Po) term describing the QCD hard processes. As shown in Ref. [1], the model parameters depend generally on the collision system (pp or Pb–Pb), but some of them are independent of the $\sqrt{s_{\mathrm{NN}}}$ or the particle type or the collision centrality.
In the present paper, a generalization of the BWTPM is developed to describe, in addition to the $p_{\textrm{T}}$-spectra, also the midrapidity elliptic flow coefficients $v_{2}$ in identical nuclei (AA) collisions. The $v_{2}$ characterizes the azimuthal asymmetry of hadron momentum distribution and depends on hadron type, $p_{\textrm{T}}$, collision system, and centrality. The asymmetry is driven mainly by an almond shape (in the plane transverse to the collision axis) of the hadronic medium (quark-gluon plasma) created in non-central collisions. Due to this shape, the in-medium pressure gradient and radial flow velocity become anisotropic resulting in the anisotropy of the hadron momentum spectra [5]. Following Ref. [6], I use the generalized BWM with a transversally anisotropic freeze-out hyper-surface of elliptical shape, instead of the isotropic circular one considered in Refs. [1, 2]. For the hadronic medium radial flow velocity I assume similar azimuthal dependence as in Refs. [6, 7]. Also, the other two therms of BWTPM are modified by including a simple dependence on $\varphi$. Having redefined so the model, the hadron $v_{2}$ can be determined as the second coefficient in the Fourier series of its invariant yield $\textrm{d}^{3}N/(p_{\textrm{T}}\textrm{d}p_{\textrm{T}}\textrm{d}y\textrm{d}\varphi)$ [6, 7, 8, 9, 10]
@@
The next objective of the present paper is to reduce strongly the number of the BWTPM parameters by parametrizing their dependence on the collision system, centrality, energy, and on the particle type and $p_{\textrm{T}}$ using simple formulae. As a measure of collision centrality, the corresponding charged-particle multiplicity density at midrapidity ($\eta$ is pseudo-rapidity)
@@
is used which serves as a good scaling variable for the comparison of hadron production mechanisms in different collision systems (see, e.g., the ALICE review paper [11]). These parametrizations increase significantly the predictive power of the model. Furthermore, I propose another simplification of the BWTPM. While in Ref. [1] it had two different functional forms for $p_{\textrm{T}}\leq 40$ GeV and $p_{\textrm{T}}>40$ GeV, it now has the same form for any $p_{\textrm{T}}$.
The values of the BWTPM parameters are obtained then from a simultaneous fit of the combined data on different hadron $p_{\textrm{T}}$-spectra and $v_{2}$ measured mostly by the ALICE experiment in Pb–Pb and Xe–Xe collisions at the LHC. In addition to the hadrons considered in Ref. [1] ($\pi^{\pm}$, $\pi^{0}$, $\mathrm{K}^{\pm}$, ${\mathrm{K}}^{0}_{\mathrm{S}}$, $\mathrm{p}$, $\mathrm{\overline{p}}$, ${\mathrm{K}}^{*0}$, $\overline{\mathrm{K}^{*0}\mkern-15.0mu}$ , $\phi$, $\Lambda$, $\overline{\Lambda}$, $\Xi^{-}$, $\overline{\Xi^{+}\mkern-12.0mu}$ , $\Omega^{-}$, $\overline{\Omega^{+}\mkern-12.0mu}$ , ${\mathrm{D}}^{0}$, $\mathrm{D}^{\pm}$, $\mathrm{J}/\psi$, and unidentified charged-particles), the $\eta$, $\rho^{0}$ mesons and $\Lambda_{c}^{+}$$\overline{\Lambda^{-}\mkern-12.0mu}_{c}$ baryons are included too. It is assumed that a particle and its antiparticle have the same $p_{\textrm{T}}$-spectrum and $v_{2}$. While in the LHC energy domain, which is the main target of present study, this assumption is perfectly valid, it could be violated at much lower energies. To test the model at such energies, I included in the simultaneous fit also some limited data on $p_{\textrm{T}}$-spectra and $v_{2}$ measured at the RHIC.
Note that the previous similar simultaneous fits of the LHC data on hadrons $p_{\textrm{T}}$-spectra and $v_{2}$ were done mainly at low $p_{\textrm{T}}$ ($p_{\textrm{T}}\lesssim 3$ GeV/c[7, 11, 12] using various generalizations of the BWM. The present model, due to its additional power-law terms, is able to describe much larger dataset in the whole available $p_{\textrm{T}}$-region.
The paper is organized as follows: Sec. II gives details of the generalized BWTPM. Sec. III presents a parametrization for the $M_{\rm ch}$ data measured in various AA collisions. Sec. IV describes parametrizations for the model parameters dependence on the AA collision characteristics and the particle type and $p_{\textrm{T}}$. In Sec. V the model free parameters are determined from the simultaneous fit of all the used data. In Sec. VI the results and predictions are discussed. Conclusions are given in the last section.
@@ II Description of BWTPM
The present model consists of three terms that generalize the similar terms of Ref. [1]. The first (BW) term corresponds to the generalized BWM [6] and two others (named Ts and Po as in Ref. [1]) include additionally the second Fourier coefficients $v_{2t}(p_{\textrm{T}})$ and $v_{2p}(p_{\textrm{T}})$, respectively, to account for the elliptic flow. Thus, for a particle $\varphi$-dependent $p_{\textrm{T}}$-spectrum in AA collisions at midrapidity, where it can be considered independent of the rapidity $y$, I propose the following model:
@@
$\rho$ is the flow rapidity with the second order anisotropy $\rho_{2}$ and profile exponent $k$, $V=\pi R_{x}R_{y}\tau_{f}$ is the volume of this medium with transverse sizes $R_{x}$ and $R_{y}$, $\tau_{f}$ and $T$ are its proper life-time and temperature, and $K_{1}$ is a modified Bessel function of the second kind. Note that $\tanh{\rho_{0}}$ is equal to $\beta_{\textrm{s}}$ used in Ref. [1]. The global normalization parameter $f_{N}$ is close to unity at LHC energies but could depend on the normalization conditions of the used data obtained in different experiments. Its energy dependence is significant at lower RHIC energies and will be specified in Sec. IV. Parameter $f_{S}$ is equal to unity (as in Ref [1]) for ordinary hadrons consisting of up and down quarks only. But for hadrons containing strange or heavier quarks it is greater than unity and quantifies the suppression of their yields depending on the hadron type, $p_{\textrm{T}}$, and collision centrality. It describes the observed enhancement of the relative yields of such hadrons with respect to ordinary ones when going from peripheral to central collisions (i.e., from small to large $M_{\rm ch}$[11]. The explicit form of $f_{S}$ is specified in Sec. IV.
@@
are explained in Refs. [6, 7]. In the present model, three parameters $R_{x},R_{y}$, and $\tau_{f}$ enter into two independent combinations via parameters $V$ and $r_{xy}\leq 1$.
The important quantities for the elliptic flow $v_{2}$ are $\rho_{2},r_{xy}$ [6, 7], $k$, and functions $v_{2t}(p_{\textrm{T}})$ and $v_{2p}(p_{\textrm{T}})$ in Eq. (II), though the $T$ and $\rho_{0}$ also influence the $p_{\textrm{T}}$-dependence of $v_{2}$ [6]. From hydrodynamic models and experimental data it follows that there is an approximately linear relation between $v_{2}$ and the eccentricity $\varepsilon_{2}$ of the collision geometry [13, 14, 15, 16], which can be defined in the present model as
@@
Moreover, in the generalized BWM [6] with $\rho_{2}=0$ one gets $v_{2}=\varepsilon_{2}/2$. I have checked by fitting the $v_{2}$ data that parameter $\rho_{2}$ as well as the functions $v_{2t}(p_{\textrm{T}})$ and $v_{2p}(p_{\textrm{T}})$ have centrality dependence similar to $\varepsilon_{2}$. For the model’s simplicity, I assume that they are proportional to $\varepsilon_{2}$. Their explicit forms will be specified in Sec. IV. All other parameters in Eq. (II) have the same meaning as in Ref. [1]. Note that parameters $f_{BW},f_{1},f_{2},c_{1}$, and $c_{2}$ are the same for any AA collision system and depend only on the particle type.
@@
Performing a similar integration in Eq. (1), one finds (see also Ref. [7])
@@
$I_{0}$ and $I_{2}$ in Eqs. (II) and (II) are modified Bessel functions of the first kind. So, one finds Eqs. (8), (II), and (II) for hadrons $p_{\textrm{T}}$-spectrum and elliptic flow, respectively. They can be used to describe the corresponding data for different identified particles. There are also similar data sets for the unidentified charged particles which are a mixture of the $\pi^{\pm},\mathrm{K}^{\pm},\mathrm{p},\mathrm{\overline{p}}$, and charged hyperons. These data have usually much larger $p_{\textrm{T}}$ range than the ones for the identified particles. Charged-particle $p_{\textrm{T}}$-spectra can be defined, assuming the same spectrum for a particle and its antiparticle, by the equations [1]
@@
Here, the functions $F_{\varphi,i}$ and $F_{i}$ are defined by Eqs. (II) and (II), respectively. The factors 2 account for the positive and negative particles and factors $p_{\textrm{T}}/m_{\textrm{T}}$ account for the change from rapidity to pseudo-rapidity at midrapidity ($d\eta/dy\approx p_{\textrm{T}}/m_{\textrm{T}}$). The last terms in Eqs. (11) and (12) describe approximately the small contribution of hyperons via the proton contribution, scaled by parameter $f_{hyp}$. Using Eqs. (11) and (1) and performing integration by $\varphi$ one can find easily the analog of Eq. (II) for the unidentified charged particles. Note that Eqs. (8), (II) and (12) for the $p_{\textrm{T}}$-spectra are similar to the corresponding ones in Ref. [1] if $f_{N}=f_{S}=1$ and $\rho_{2}=0$.
@@
It was checked that the presented generalized BWTPM is able to fit the LHC data on midrapidity hadron $p_{\textrm{T}}$-spectra and $v_{2}$ in Pb–Pb and Xe–Xe collisions with a good accuracy, similar to that achieved in Ref. [1]. However, as in Ref. [1], the number of the model free parameters is too big, which reduces the model’s predictive power. It is mostly due to the ten centrality-dependent parameters $V,V_{1},V_{2},T,k,e_{2},n,f_{S},r_{xy},$ and $\rho_{0}$ which generally can depend also on $p_{\textrm{T}}$, the collision system and $\sqrt{s_{\mathrm{NN}}}$ ($f_{S}$ depends on particle species too). A typical number of ALICE centrality classes for Pb–Pb data at $\sqrt{s_{\mathrm{NN}}}=$ 2.76 TeV [17] and 5.02 TeV [18] as well as for Xe–Xe data at $\sqrt{s_{\mathrm{NN}}}=$ 5.44 TeV  [19] is about 10 in each case. So, to describe these combined data, one needs to fit $10(10+10+10)=300$ independent values for the mentioned ten parameters. Taking into account also the RHIC data included in the present study, this number of independent values becomes more than 600. To reduce the number of free parameters strongly, these ten (and other) parameters are parametrized using the charged-particle multiplicities $M_{\rm ch}$ which depend on the collision system, centrality and $\sqrt{s_{\mathrm{NN}}}$, and are well known experimentally for various AA collisions  [17, 18, 19, 20, 21, 22]. The key assumption of the present paper is that all the centrality dependence in BWTPM can be expressed via $M_{\rm ch}$ only. It is motivated by the experimental observation that any hadron species has similar $p_{\textrm{T}}$-spectra in different AA collisions with the same $M_{\rm ch}$ and $\sqrt{s_{\mathrm{NN}}}$ (see an example of such a similarity for ${\mathrm{K}}^{*0}$ meson $p_{\textrm{T}}$-spectra in Ref. [23]). To have a more general and predictive approach, I will use not the experimental values of $M_{\rm ch}$ (which can be missing for some centralities), but their accurate parametrization, valid for any AA collision centrality and energy. This parametrization is presented in the next section.
III parametrization of $M_{\rm ch}$
@@
The mean value of charged-particle multiplicity density at midrapidity $M_{\rm ch}$ (defined in Eq. (2)) is one of the main characteristics of nucleus-nucleus collisions. Its dependence on the collision system, centrality and $\sqrt{s_{\mathrm{NN}}}$ was measured for different AA collisions at LHC [17, 18, 19, 20] and RHIC [21, 22] energies. Centrality is commonly expressed in percentiles or fractions $x_{c}$ of the total nucleus-nucleus cross section [11] (centrality in % is equal to $100x_{c}$). Here I propose a simple parametrization for the $M_{\rm ch}$ data [17, 18, 19, 20, 21, 22] versus atomic mass number A, collision centrality fraction $x_{c}$ and energy $\sqrt{s_{\mathrm{NN}}}$. It relates $M_{\rm ch}$ with the similar quantity $M^{\rm pp}_{\rm ch}\equiv\langle\textrm{d}N^{\rm pp}_{\rm ch}/\textrm{d}\eta\rangle_{\eta=0}$ of the inelastic pp collisions at energy $\sqrt{s}=\sqrt{s_{\mathrm{NN}}}$. For $M^{\rm pp}_{\rm ch}$ I used the data measured at the LHC [24, 25, 26] as well as at lower energies [21, 27] and parametrized their energy dependence. The $M_{\rm ch}$ and $M^{\rm pp}_{\rm ch}$ data were obtained by averaging the respective multiplicity distribution on pseudo-rapidity in the range $|\eta|<0.5$ [17, 18, 19, 20, 22, 24, 25, 26] or $|\eta|<1$ [21, 27]. Within the data uncertainties both ranges give similar results [26]. Parameters of the parametrization are obtained via a simultaneous fit of the $M_{\rm ch}$ and $M^{\rm pp}_{\rm ch}$ data using the combined fit technique of the ROOT framework [28], which gives $\chi^{2}/NDF=54.1/155$. The fit results are shown in Figures 1 and 2 and the parametrization has the form
@@
where $\alpha=\log(\frac{17.4}{\sqrt{s_{\mathrm{NN}}}}),\beta=(\frac{0.0027}{\sqrt{s_{\mathrm{NN}}}})^{0.054},\delta=\textrm{A}^{0.063},\gamma=1+1.22x_{c}^{0.47},y=\textrm{A}(1-x_{c})$, and $e_{0}=0.005$ TeV (everywhere in the present paper, if it is not specified otherwise, $\sqrt{s_{\mathrm{NN}}}$ is defined in units of TeV). The bottom panel of Fig. 1 demonstrates that Eq. (13) describes most of the data points with better than 10% accuracy. However, it underestimates the ALICE recent data for the most peripheral Pb–Pb collisions at 5.36 TeV [20]. This could be due to the new centrality calibration [20], since all other ALICE data are described well. One can expect that Eq. (13) would give accurate values of $M_{\rm ch}$ also for new AA collisions. The result for oxygen–oxygen (OO) collisions, carried out at the LHC in July 2025, is also shown in Fig. 1. Note that for $\textrm{A}=1$ or for the most peripheral collisions, when $x_{c}\to 1$, the right-hand side of Eq. (13) equals unity. For most central collisions, when $x_{c}\to 0$, one obtains a very simple formula for the maximal values of $M_{\rm ch}$ versus A and $\sqrt{s_{\mathrm{NN}}}$:
$\frac{M_{\rm ch}^{max}}{M^{\rm pp}_{\rm ch}}=1+(\textrm{A}-1)^{1+0.00013\textrm{A}}\bigl(\frac{\sqrt{s_{\mathrm{NN}}}}{0.0027}\bigr)^{0.1134}.$ (15)
For example, $M_{\rm ch}^{max}=$ 2404 (2457) for Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}=$5.02 (5.36) TeV. It is worth noting that Eq. (13) should be used with caution since it gives the same $M_{\rm ch}$ for collisions of isobars of the same A [29].